OpenStudy (anonymous):

If the function rule is x - 1/3 = y and x = -3, what is the value of y?

OpenStudy (anonymous):

-3 1/3

OpenStudy (anonymous):

x - 1/3 = y, where x = -3 -3 -1/3 = y (-9/3) - 1/3 = y (multiplying numerator and denomiator by 3) -10/3 = y So y = -10/3 (or) -3 1/3

OpenStudy (anonymous):

-3 1/3

OpenStudy (anonymous):

x-1/3=y (x=-3) -3-1/3=y -4/3=y